Based on the frequency distribution above, find the relative frequency for the class with lower class limit 27.

AgesNumber of students
15-189
19-222
23-262
27-308
31-3410
35-3810


Based on the frequency distribution above, find the relative frequency for the class with lower class limit 27.


Relative Frequency = _______ %
Answer #1

Number of students in class with lower class limit 27, that is, 27- 30 = 8

Relative frequency = 8/(9+2+2+8+10+10) = 8/41 = 0.1951

In percentage = 19.51 = 19.5%

Hope this helps

Please up vote.